OpenStudy (khally92):

No 5 please. calculus 3

OpenStudy (khally92):

@Loser66 Help with no 5 please. just wanna double check and make sure I am doing the right thing. Thank you..

OpenStudy (anonymous):

Where is the question?

OpenStudy (khally92):

@Lovelarap No 5 please. thank you

OpenStudy (khally92):

Thank you.

OpenStudy (khally92):

I got dz/dt=-3(df/dx)+2(df/dy)

OpenStudy (anonymous):

Sorry, I can't really help you because i'm still on pre-calc.

OpenStudy (khally92):

Oh Ok, No problem lets wait for @Loser66 .

OpenStudy (khally92):

@Astrophysics @ShadowLegendX No 5 please.

OpenStudy (loser66):

show your work first

OpenStudy (khally92):

Ok \[x=2-3t \] \[y=3+2t\] \[dx/dt= -3\] \[dy/dt=2\] Therefore \[dz/dt=(df/dx)(dx/dt)+(df/dy)(dy/dt)\] \[dz/dt=-3(df/dx)+2(df/dy)\]

OpenStudy (khally92):

@Loser66 .

OpenStudy (khally92):